Mercy Corporation"s bond have 15years to maturity, an 8.75% coupon paid semi- annually, and a $1,000 par value.The bond has a 6.5% nominal yield to maturity, but it can be cslled in 6years at a price of $1,050. What is the bond's nominal yield to call?
Answer #1

Par Value =1000
Number of Periods =15*2 =30.
semi annual coupon =8.75%*1000/2 =43.75
Price of Bond =PV of Coupons+PV of Par Value =43.75*((1-(1+3.25%)^-30)/3.25%)+1000/(1+3.25%)^30 =1213.55

Number of Periods of call =6*2 =12
Call Price =1050
YTC using financial calculator
N=12;PMT=43.75;PV=-1213.55;FV=1050;CPT I/Y =2.316%
YTC =2*2.316% =4.63%